The time it takes a mechanic to complete an oil change is 7 minutes for each change and 10 minutes to set up the station for the day. Write a function where x represents the number of cars that get an oil change. How long does it take to complete 50 oil changes?
Thanks
Christina

Answer by stanbon(75887)

The time it takes a mechanic to complete an oil change is 7 minutes for each change and 10 minutes to set up the station for the day. Write a function where x represents the number of cars that get an oil change. How long does it take to complete 50 oil changes?
----
Time for 50 changes:: 10 + 50*7 = 360 minutes = 360/60 = 6 hrs
----------------
Cheers,
Stan H.
---------